# A note on Lie and Jordan structures of Leavitt path algebras

### v1

@@ 4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
    If $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable, then it is Lie solvable of index $l\leq 3$. More precisely, if $|I|$ is finite and ${\rm char}(K)\neq 2$ ((((resp., ${\rm char}(K)=2$)))), then $l=1$ ((((resp., $l=2$)))). If $|I|$ is infinite and ${\rm char}(K)\neq 2$ ((((reps., ${\rm char}(K)=2$)))), then $l=2$ ((((resp., $l=3$)))). Additionally, $\mathbf{K}_{L_{K}(E_{4})}$ is Lie nilpotent if and only if $|I|$ is finite and ${\rm char}(K)\neq 2$, if and only if $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]=1$.
@@
    $\mathbf{K}_{L_{K}(E_{5})}$ is Lie solvable ((((of index $l\leq 4$)))) if and only if ${\rm char}(K)=2$. More precisely, if $|J|$ is finite ((((resp., infinite)))), then $l=3$ ((((resp., $l=4$)))). Additionally, $\mathbf{K}_{L_{K}(E_{5})}$ is not Lie nilpotent.
@@
    $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable ((((of index $l\leq 4$)))) if and only if ${\rm char}(K)=2$. More precisely, if both $|I|$ and $|J|$ are finite ((((resp., either $|I|$ or $|J|$ is infinite)))), then $l=3$ ((((resp., $l=4$)))). Additionally, $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent.
@@
Moreover, according to Lemma 4.1, the set $\{u_{i},e_{i},e_{i}^{*},e_{i}e_{i}^{*}\}$ is linearly independent over $K$. Thus, it is can be checked that $I_{i}\cong\mathbb{M}_{2}(K)$ via the mapping $u_{i}\mapsto E_{11}$, $e_{i}e_{i}^{*}\mapsto E_{22}$, $e_{i}\mapsto E_{12}$ and $e_{i}^{*}\mapsto E_{21}$. Moreover, we have $\sum_{i=1}^{n}I_{i}=\bigoplus_{i=1}^{n}I_{i}$. Put $M=\bigoplus_{i=1}^{n}I_{i}$. Then, according to Lemma 4.3 and Corollary 3.2, we conclude that $\mathbf{K}_{M}=\bigoplus_{i=1}^{n}\mathbf{K}_{I_{i}}$, and that $\mathbf{K}_{M}$ is Lie solvable. Moreover, in view of Corollary 3.2, if ${\rm char}(K)\neq 2$, then $[\mathbf{K}_{M},\mathbf{K}_{M}]=0$; and, if ${\rm char}(K)=2$ then $\mathbf{K}_{M}$ is Lie solvable of index $2$ but not Lie nilpotent. Now, we consider two possible cases:
@@
$[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq[L_{K}(E_{4}),L_{K}(E_{4})]=[M,M]\subseteq M.$
As $M$ is Lie solvable, we conclude that $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$, and hence $\mathbf{K}_{L_{K}(E_{4})}$, is Lie solvable. To determine the solvable index of $\mathbf{K}_{L_{K}(E_{4})}$, we consider the following subcases:
@@
Subcase 2.2. ${\rm char}(K)=2$. As $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq\mathbf{K}_{M}$ which is Lie solvable of index $2$, it follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index at most $3$. Because $v\in\mathbf{K}_{L_{K}(E_{4})}\backslash\mathbf{K}_{M}$, we conclude that $\mathbf{K}_{L_{K}(E_{4})}\neq[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$. It follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index $3$. On the other hand, as $\mathbf{K}_{I_{i}}$ is not Lie nilpotent and $\mathbf{K}_{I_{i}}\subseteq\mathbf{K}_{L_{K}(E_{4})}$ for all $i$, we obtain that $\mathbf{K}_{L_{K}(E_{4})}$ is not Lie nilpotent. The proof of assertion (1) is complete.
@@
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 4$)))) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Specially, the index is $4$ precisely when $E$ contains either $E_{5}$ or $E_{6}$ with infinite $|I|$ or $|J|$.
@@
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 2$)))) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$. Specially, the index is $2$ precisely when $E$ contains $E_{4}$ with infinite $|I|$.
@@
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 3$)))) if and only if $E$ is a disjoint union graphs $E_{1},\dots,E_{6}$ with finite $|I|$ and $|J|$. In this case,
@@
    for suitable ((((possibly infinite)))) indexing sets $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$. Moreover, the index is $3$ precisely when $E$ contains $E_{3}$, $E_{5}$ or $E_{6}$.
@@
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 2$)))) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$ with finite $|I|$. In this case, we have
@@
    for suitable ((((possibly infinite)))) indexing sets $J_{1}$, $J_{2}$, $J_{3}$. Moreover, the index is $2$ precisely when $E$ contains $E_{3}$.
@@
In particular, $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $[\mathbf{K}_{L_{K}(E)},\mathbf{K}_{L_{K}(E)}]=0$.
@@
    If ${\rm char}(K)=2$, then $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable ((((resp., nilpotent)))) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$ ((((resp., $E_{1}$ and $E_{2}$)))).
@@
Funding. This research is funded by the Vietnam Ministry of Education and Training under grant number B2024-CTT-02.
Conflict of Interest. The authors have no conflict of interest to declare that are relevant to this article.
Data availability. The authors declare that the data supporting the findings of this study are available within the paper.

### v2

@@ 4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
    If $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable, then it is Lie solvable of index $l\leq 3$. More precisely, if $|I|$ is finite and ${\rm char}(K)\neq 2$ ((resp., ${\rm char}(K)=2$)), then $l=1$ ((resp., $l=2$)). If $|I|$ is infinite and ${\rm char}(K)\neq 2$ ((reps., ${\rm char}(K)=2$)), then $l=2$ ((resp., $l=3$)). Additionally, $\mathbf{K}_{L_{K}(E_{4})}$ is Lie nilpotent if and only if $|I|$ is finite and ${\rm char}(K)\neq 2$, if and only if $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]=1$.
@@
    $\mathbf{K}_{L_{K}(E_{5})}$ is Lie solvable ((of index $l\leq 4$)) if and only if ${\rm char}(K)=2$. More precisely, if $|J|$ is finite ((resp., infinite)), then $l=3$ ((resp., $l=4$)). Additionally, $\mathbf{K}_{L_{K}(E_{5})}$ is not Lie nilpotent.
@@
    $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable ((of index $l\leq 4$)) if and only if ${\rm char}(K)=2$. More precisely, if both $|I|$ and $|J|$ are finite ((resp., either $|I|$ or $|J|$ is infinite)), then $l=3$ ((resp., $l=4$)). Additionally, $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent.
@@
Moreover, according to Lemma 4.1, the set $\{u_{i},e_{i},e_{i}^{*},e_{i}e_{i}^{*}\}$ is linearly independent over $K$. Thus, it is can be checked that $I_{i}\cong\mathbb{M}_{2}(K)$ via the mapping $u_{i}\mapsto E_{11}$, $e_{i}e_{i}^{*}\mapsto E_{22}$, $e_{i}\mapsto E_{12}$ and $e_{i}^{*}\mapsto E_{21}$. Moreover, we have $\sum_{i=1}^{n}I_{i}=\bigoplus_{i=1}^{n}I_{i}$. Put $M=\bigoplus_{i=1}^{n}I_{i}$. Then, according to Lemma 4.3 and Corollary 3.2, we conclude that $\mathbf{K}_{M}=\bigoplus_{i=1}^{n}\mathbf{K}_{I_{i}}$, and that $\mathbf{K}_{M}$ is Lie solvable. Moreover, in view of Corollary 3.2, if ${\rm char}(K)\neq 2$, then $\mathbf{K}_{M},\mathbf{K}_{M}]=0$; and, if ${\rm char}(K)=2$ then $\mathbf{K}_{M}$ is Lie solvable of index $2$ but not Lie nilpotent. Now, we consider two possible cases:
@@
$\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq[L_{K}(E_{4}),L_{K}(E_{4})]=[M,M]\subseteq M.$
As $M$ is Lie solvable, we conclude that $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$, and hence $\mathbf{K}_{L_{K}(E_{4})}$, is Lie solvable. To determine the solvable index of $\mathbf{K}_{L_{K}(E_{4})}$, we consider the following subcases:
@@
Subcase 2.2. ${\rm char}(K)=2$. As $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq\mathbf{K}_{M}$ which is Lie solvable of index $2$, it follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index at most $3$. Because $v\in\mathbf{K}_{L_{K}(E_{4})}\backslash\mathbf{K}_{M}$, we conclude that $\mathbf{K}_{L_{K}(E_{4})}\neq[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$. It follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index $3$. On the other hand, as $\mathbf{K}_{I_{i}}$ is not Lie nilpotent and $\mathbf{K}_{I_{i}}\subseteq\mathbf{K}_{L_{K}(E_{4})}$ for all $i$, we obtain that $\mathbf{K}_{L_{K}(E_{4})}$ is not Lie nilpotent. The proof of assertion (1) is complete.
@@
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 4$)) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Specially, the index is $4$ precisely when $E$ contains either $E_{5}$ or $E_{6}$ with infinite $|I|$ or $|J|$.
@@
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 2$)) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$. Specially, the index is $2$ precisely when $E$ contains $E_{4}$ with infinite $|I|$.
@@
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 3$)) if and only if $E$ is a disjoint union graphs $E_{1},\dots,E_{6}$ with finite $|I|$ and $|J|$. In this case,
@@
    for suitable ((possibly infinite)) indexing sets $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$. Moreover, the index is $3$ precisely when $E$ contains $E_{3}$, $E_{5}$ or $E_{6}$.
@@
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 2$)) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$ with finite $|I|$. In this case, we have
@@
    for suitable ((possibly infinite)) indexing sets $J_{1}$, $J_{2}$, $J_{3}$. Moreover, the index is $2$ precisely when $E$ contains $E_{3}$.
@@
In particular, $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $\mathbf{K}_{L_{K}(E)},\mathbf{K}_{L_{K}(E)}]=0$.
@@
    If ${\rm char}(K)=2$, then $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable ((resp., nilpotent)) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$ ((resp., $E_{1}$ and $E_{2}$)).
@@
Funding. This work is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under Grant No. 101.04-2025.41.